I need the steps and answer to this :)
|x - 2| + 3 = 10
: I need the steps and answer to this :)
|x - 2| + 3 = 10

Answer by oscargut(667)

|x - 2| + 3 = 10 then
|x - 2| = 7
then x-2=7 or x-2=-7
then x=9 or x=-5
Answer: x=-5,x=9